OpenStudy (anonymous):

solve the equation for M F=-GMn/D^2

OpenStudy (anonymous):

FD^2=-GMn FD^2/-Mn=G

OpenStudy (saifoo.khan):

Cross multiply, \[D^2 F = GMn\] Divide G and n from both sides. \[\frac{D^2 F}{Gn}= M\]

OpenStudy (anonymous):

multiply both sides by d^2 then divide both sides by -Mn

OpenStudy (anonymous):

g=d^2/-mn

OpenStudy (anonymous):

oh my bad i solved for G sorry! saifoo is right